OpenStudy (anonymous):

The area of a rectangle can be found by multiplying the length by the width. A rectangular postage stamp has a length of nine-tenths inch and a width of five-sixths inch. What is the area of the stamp, in square inches? three-fourths square inches one-seventeenth square inches one-half square inches one and one-fourth square inches

OpenStudy (anonymous):

A = L x W A = 9/10 x 5/6 A = 45/60 which reduces to 3/4 sq inches

OpenStudy (anonymous):

than you

OpenStudy (anonymous):

thank you